OpenStudy (anonymous):

@StudyGurl14 17 - 5m = 20 + 3m This one i'm really confused on because i do not know if it is -5m

OpenStudy (studygurl14):

what do you mean "do not know if it is -5m" ?

OpenStudy (anonymous):

Because it says 17 - 5m i do not know if it meant as a -5m or 5m

OpenStudy (studygurl14):

It means "17 minus 5m" However, it can also be read as "17 plus negative 5m"

OpenStudy (studygurl14):

First step. Let's move the 3m to the left side

OpenStudy (studygurl14):

What is the opposite of 3m?

OpenStudy (anonymous):

-3m

OpenStudy (studygurl14):

correct. so you would subtract 3m from both sides of the equation 17 - 5m = 20 + 3m - 3m - 3m ____________________ 17 - 8m = 20 + 0 17 - 8m = 20 What now?

OpenStudy (anonymous):

do the opposite of subtraction

OpenStudy (studygurl14):

umm...not really. You want to move the 17 to the other side. How do you do that?

OpenStudy (anonymous):

by adding 17 to 20

OpenStudy (studygurl14):

almost. Since 17 is positive, you want to subtract 17 from 20, not add

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Which is 3 right?

OpenStudy (studygurl14):

right

OpenStudy (studygurl14):

so you end up with -8m = 3

OpenStudy (studygurl14):

what next?

OpenStudy (anonymous):

m = -0.375

OpenStudy (studygurl14):

right

OpenStudy (studygurl14):

also -3/8

OpenStudy (anonymous):

Thank you

OpenStudy (studygurl14):

yw